# Two-dimensional fluids via matrix hydrodynamics

### v2

@@ 4. Coadjoint orbits and their closure
However, while solutions with smooth initial data remain smooth, the $C^{k}$-norms for $k\geq 1$ typically grow fast as the vorticity level-curves quickly become extremely entangled. This mechanism results in a Reeb graph that is well-defined but futile to track at moderate scales. It is therefore natural to study the closure of coadjoint orbits in a weaker topology, for instance $L^{\infty}$ as Yudovich’ existence theory suggests. In doing so, the measured Reeb graph is not stable in $L^{\infty}$ since it is based on critical values and Morse functions that require $C^{2}$ topology (see Fig. 3). Instead, what persists is the level-set measure $\lambda_{\omega}$, obtained by “flattening” the Reeb graph as follows. The mapping $C^{\infty}(\mathbb{R})\ni f\mapsto C_{f}(\omega)$ is a distribution with compact support, and $\lambda_{\omega}$ is the corresponding positive Radon measure on $\mathbb{R}$ (see Fig. 2b). The $L^{\infty}$ closure of a coadjoint orbit then fulfills $\overline{\mathcal{O}(\omega_{0})}\subseteq\{\omega\in L^{\infty}(\mathbb{S}^{2})\mid\lambda_{\omega}=\lambda_{\omega_{0}}\}$.
@@
$C_{z^{m}}(W_{N})=\frac{(-i)^{m}}{N}\operatorname{tr}(W_{N}^{m})=\frac{(-i)^{m}}{N}\operatorname{tr}(W_{N}W_{N}^{m-1})=$
@@
Since the scaled Frobenius inner product $-\frac{1}{N}\operatorname{tr}(AB)$ on $\mathfrak{u}(N)$ corresponds to the $L^{2}$ inner product, i.e., $\langle T_{N}^{*}A,T_{N}^{*}B\rangle_{L^{2}}=-\frac{2\pi}{N}\operatorname{tr}(AB)$, we have
$\lim_{N\to\infty}\frac{(-i)^{m}}{N}\operatorname{tr}(W_{N},T_{N}(\omega^{m-1}))=\frac{1}{2\pi}\langle\omega,\omega^{m-1}\rangle_{L^{2}}=C_{z^{m}}(\omega).$
@@
which follows from Lemma 7 below ∎
@@
The result is true for $m=1$. Assume it for $m-1$. With $W_{N}=T_{N}(\omega)$ we have
@@
That the first term vanishes as $N\to\infty$ follows from the assumption. That the second term vanishes follows from product estimates in quantization theory [3, prop. 3.12]. ∎
[…]
@@
The semi-norm $\lvert\cdot\rvert_{N}$ can “see” structures down to the length scale $\hbar_{N}$. Furthermore, each length scale $\hbar_{N}$ gives an associated truncated level-set measure $\lambda_{T_{N}(\omega)}^{N}$. To study how it varies, for fixed $N$ as $t\to\infty$, gives information about vortex mixing relative to the scale $\hbar_{N}$ and a connection to random matrix theory (see Fig 4).
@@
$\lvert\langle\omega,f\rangle\rvert=\lvert\langle\omega,T_{N}^{*}W\rangle\rvert=\frac{2\pi}{N}\lvert\operatorname{tr}(WT_{N}\omega)\rvert\leq\frac{2\pi}{N}\lvert\omega\rvert_{N}\lVert W\rVert_{1}.$
@@
Thus, a plausible relaxation of Euler’s equations (5) to the weak-* topology is to consider the Euler–Zeitlin equation (7) simultaneously for all $N\geq N_{0}$. Via the mappings $T_{N-1}\circ T^{*}_{N}\colon\mathfrak{u}(N)\to\mathfrak{u}(N-1)$ one can consider an inverse limit $\mathfrak{g}=\underleftarrow{\lim}\,\mathfrak{u}(N)$. But this construction is too rigid, since $\mathfrak{g}$ is not invariant under the simultaneous Euler–Zeitlin flow. An open problem is to relax the inverse limit, for example based on the equivalence $\{W^{\prime}_{i}\}_{i=N_{0}}^{\infty}\sim\{W_{i}\}_{i=N_{0}}^{\infty}$ if $\lim_{i\to\infty}\lVert W^{\prime}_{i}-W_{i}\rVert_{\infty}=0$, and then prove that the simultaneous Euler–Zeitlin flow is well-defined on equivalence classes. For such a construction, global existence is built-in (albeit not uniqueness).
@@ 5. Canonical splitting and mixing operators
Stationary solutions of the Euler–Zeitlin equations (7) are characterized by the relation $[P,W]=0$. The stabilizer of $P$ is the subspace $\mathrm{stab}_{P}=\{W\in\mathfrak{su}(N)\mid[P,W]=0\}$. For a state $W$ and $P=\Delta_{N}^{-1}W$, the distance of $W$ to $\mathrm{stab}_{P}$ thus measures how far from stationary the state is.
Let $\Pi_{P}$ denote the $L^{2}$ projection onto $\mathrm{stab}_{P}$. It induces a canonical splitting $W=W_{s}+W_{r}$ where $W_{s}=\Pi_{P}W$. Explicitly, if $E\in\mathrm{U}(N)$ is an eigen basis for $P$, then $\Pi_{P}(W)=E^{*}\mbox{diag}(EWE^{*})E$, where diag is the orthogonal projection onto the space of diagonal matrices. From this formula, we can express the Euler–Zeitlin equations in the variables $W_{s}$ and $W_{r}$:
@@
where $B$ is the unique element in the $L^{2}$ complement $\mathrm{stab}_{P}^{\perp}$ such that $[B,P]=\Pi_{P}^{\perp}\Delta^{-1}[P,W]$. By construction, the decomposition is orthogonal in $L^{2}$. Curiously, it also fulfills a “reversed” orthogonally condition with respect to the energy norm $\lVert W\rVert_{E}^{2}\coloneqq\langle W,\Delta_{N}^{-1}W\rangle_{2}$.
@@
Why is the canonical splitting interesting? The theory of statistical mechanics, applied to Euler’s equations on $\mathbb{S}^{2}$, predicts a long-time behavior of small scale fluctuations about a large scale, steady state [15]. Thus, we expect $W_{r}=W-W_{s}$ to decrease [^3] But due to Poincaré recurrence only in a statistical sense: see zeitlin-reversibility-blog [29]. with time in the energy norm, which is in-sensitive to small scales. Indeed, in numerical simulations the component $W_{s}$ favors evolution into large scales, while $W_{r}$ favors small scales (see Fig. 5 and [32, fig. 3b]). This canonical separation of scales is a perspective on Kraichnan’s [26] backward energy and forward enstrophy cascades.
@@
The analog of $\Pi_{P}$ is the operator $\Pi_{\psi}$ for $L^{2}$ projection onto $\mbox{stab}_{\psi}=\overline{\operatorname{ker}\{\psi,\cdot\}}^{L^{2}}$, formally given by averaging along the level-sets of $\psi$. For a generic $\psi\in C^{1}(\mathbb{S}^{2})$, it is well-defined as a norm-1 operator $\Pi_{\psi}\colon L^{p}(\mathbb{S}^{2})\to L^{p}(\mathbb{S}^{2})$ [32, prop. 5.4].
Lemma 10 ([32]).
@@
Let $\{\omega_{n}\}$ be a uniformly bounded sequence in $L^{\infty}$. Then $\|\omega_{n}-\overline{\omega}\|_{-1}\rightarrow 0$ if and only if $\omega_{n}\overset{\ast}{\rightharpoonup}\overline{\omega}$ weakly-*- ∗ in $L^{\infty}$ if and only if $\omega_{n}\rightharpoonup\overline{\omega}$ weakly in $L^{2}$, for $n\rightarrow\infty$.
@@
Weak-*- ∗ convergence in $L^{\infty}$ clearly implies weak convergence in $L^{2}$, being $L^{2}\subset L^{1}$, for bounded domains.
On the other hand, if $\omega_{n}\rightharpoonup\overline{\omega}$ weakly in $L^{2}$, then being $L^{2}$ compactly embedded in $H^{-1}$ by the Rellich–Kondrachov theorem, we have that $\|\omega_{n}-\overline{\omega}\|_{-1}\rightarrow 0$.
Finally, it remains to prove that convergence in $H^{-1}$ implies convergence in the weak-*- ∗ topology. Let $\phi\in C^{1}$, then we have that
[…]
@@
[…]
@@
The thesis follows immediately from Lemma 12 and from the uniform bound
$\|\omega_{r}(t_{n})\|_{\infty}\leq\|\omega(0)\|_{\infty},$
for any $n\geq 1$. ∎
Hence, the states $\omega_{s}$, observed in our long-time numerical simulations, are natural candidates for the weak-* vorticity limits. We notice that minimal flows seem too restrictive as candidates for long-time asymptotics on $\mathbb{S}^{2}$; essentially, they allow only two vortex condensates, or the shape of similar signed condensates would have to balance perfectly, which is physically unlikely and numerically unseen. Instead, it is common to see branching in the functional dependence between vorticity and stream function [32, fig. 7 and 8].
@@
where $M$ is the angular momentum. We then expect, up to a global rotation, weak-* limits of the Euler equations as
$\Pi_{\psi(t_{n})}\omega(t_{n})\overset{\ast}{\rightharpoonup}\overline{\omega},\qquad n\to\infty.$

### v3

@@ 4. Coadjoint orbits and their closure
However, while solutions with smooth initial data remain smooth, the $C^{k}$-norms for $k\geq 1$ typically grow fast as the vorticity level-curves quickly become extremely entangled. This mechanism results in a Reeb graph that is well-defined but futile to track at moderate scales. It is therefore natural to study the closure of coadjoint orbits in a weaker topology, for instance $L^{\infty}$ as Yudovich’ existence theory suggests. In doing so, the Reeb graph is not stable in $L^{\infty}$ since it is based on critical values and Morse functions that require $C^{2}$ topology (see Fig. 3). Instead, what persists is the level-set measure $\lambda_{\omega}$, obtained by “flattening” the Reeb graph as follows. Recall the Casimir invariants $C_{f}(\omega)$. The mapping $C^{\infty}(\mathbb{R})\ni f\mapsto C_{f}(\omega)$ is a distribution with compact support, and $\lambda_{\omega}$ is the corresponding positive Radon measure on $\mathbb{R}$ (see Fig. 2b). The $L^{\infty}$ closure of a coadjoint orbit then fulfills $\overline{\mathcal{O}(\omega_{0})}\subseteq\{\omega\in L^{\infty}(\mathbb{S}^{2})\mid\lambda_{\omega}=\lambda_{\omega_{0}}\}$. (The latter set is characterized in Sec. 4.1 below.)
@@
$\frac{1}{4\pi}C^{N}_{z^{m}}(W_{N})=\frac{(-i)^{m}}{N}\operatorname{tr}(W_{N}^{m})=\frac{(-i)^{m}}{N}\operatorname{tr}(W_{N}W_{N}^{m-1})=$
@@
Since the scaled Frobenius inner product on $\mathfrak{u}(N)$ corresponds to the $L^{2}$ inner product, i.e., $\langle T_{N}^{*}A,T_{N}^{*}B\rangle_{L^{2}}=-\frac{4\pi}{N}\operatorname{tr}(AB)$, we have
$\lim_{N\to\infty}\frac{(-i)^{m}}{N}\operatorname{tr}(W_{N},T_{N}(\omega^{m-1}))=\frac{1}{4\pi}\langle\omega,\omega^{m-1}\rangle_{L^{2}}=\frac{1}{4\pi}C_{z^{m}}(\omega).$
@@
which follows from Lemma 7 below. ∎
@@
The result is true for $m=1$. Assume it for m1m-1. With $W_{N}=T_{N}(\omega)$ we have
@@
That the first term vanishes as $N\to\infty$ follows from the assumption. That the second term vanishes follows from product estimates in Berezin-Toeplitz quantization theory [8, prop. 3.12] and its equivalence to quantization via representations, as described in Appendix B. ∎
The level-set measure $\lambda_{\omega}$ is invariant under the Euler flow (6). Likewise, $\lambda^{N}_{W}$ is invariant under the Euler–Zeitlin flow (9). In particular, for any $f\in C^{\infty}(\mathbb{R})$ the function $C^{N}_{f}(W)=\int_{\mathbb{R}}f\,d\lambda^{N}_{W}$ is a Casimir for $\mathfrak{su}(N)$. A direct corollary of Theorem 6 is that $C^{N}_{f}(T_{N}(\omega))\to C_{f}(\omega)$ as $N\to\infty$. In this sense, the Euler–Zeitlin equation nearly conserves all Casimirs of the vorticity equation.
@@
The semi-norm $\lvert\cdot\rvert_{N}$ can “see” structures down to the length scale $\hbar_{N}$. Furthermore, each length scale $\hbar_{N}$ gives an associated truncated level-set measure $\lambda_{T_{N}(\omega)}^{N}$. To study how it varies, for fixed $N$ as $t\to\infty$, gives information about vortex mixing relative to the scale $\hbar_{N}$ and a connection to random matrix theory (see Fig. 4).
@@
$\lvert\langle\omega,f\rangle\rvert=\lvert\langle\omega,T_{N}^{*}W\rangle\rvert=\frac{4\pi}{N}\lvert\operatorname{tr}(WT_{N}\omega)\rvert\leq\frac{4\pi}{N}\lvert\omega\rvert_{N}\lVert W\rVert_{1}.$
@@
Remark 5.
Alternatively, one can view Lemma 8 as a special case of the general principle that two Hausdorff topologies on a set are equivalent if the first is compact and the second is weaker than the first.
@@
In consequence, a plausible relaxation of Euler’s equations (6) to the weak-* topology is to consider the Euler–Zeitlin equation (9) simultaneously for all $N\geq N_{0}$. Via the mappings $T_{N-1}\circ T^{*}_{N}\colon\mathfrak{u}(N)\to\mathfrak{u}(N-1)$ one can consider an inverse limit $\mathfrak{g}=\underleftarrow{\lim}\,\mathfrak{u}(N)$. But this construction is too rigid, since $\mathfrak{g}$ is not invariant under the simultaneous Euler–Zeitlin flow. A plausible direction is to relax the inverse limit, for example based on the equivalence $\{W^{\prime}_{i}\}_{i=N_{0}}^{\infty}\sim\{W_{i}\}_{i=N_{0}}^{\infty}$ if $\lim_{i\to\infty}\lVert W^{\prime}_{i}-W_{i}\rVert_{\infty}=0$, and attempt to prove that the simultaneous Euler–Zeitlin flow is well-defined on equivalence classes. For such a construction, global existence is built-in (albeit not uniqueness).
4.1. Schur-Horn-Kostant convexity
In their paper on Schur-Horn-Kostant convexity for the area-preserving diffeomorphism group of the annulus, Bloch, Flaschka, and Ratiu [5] (BFR) write: “Diffeomorphism groups are huge, infinite-dimensional Lie groups, but in some respects they can be strikingly similar to finite-dimensional semisimple groups.” Matrix hydrodynamics can be viewed as an actualization of this statement.
Consider the annulus (or cylinder)
$\mathbb{A}=\mathbb{S}^{1}\times[-1,1]=\{(\exp(\mathrm{i}\phi),z)\mid 0\leq\phi<2\pi,-1\leq z\leq 1\}$
equipped with the area-form $\nu=d\phi\wedge dz$. Let $\mathrm{Diff}_{\nu}(\mathbb{A})$ denote the infinite-dimensional Lie group [^3] More precisely, $\mathrm{Diff}_{\nu}(\mathbb{A})$ is an infinite-dimensional Lie group in the category of Fréchet manifolds (cf. Hamilton [24, Thm. 2.5.3]). of area-preserving diffeomorphisms. Its Lie algebra $\mathfrak{a}$ can be identified with the space of smooth Hamiltonian functions on $\mathbb{A}$, modulo constants, that are constant on the edges z=1z=-1 and $z=1$ (corresponding to tangential boundary conditions), equipped with the Poisson bracket induced by $\nu$. The point of BFR is to think morally of the maps $(\phi,z)\mapsto(\phi+\alpha(z),z)$ as the maximal Abelian subgroup $\mathrm{D}\subset\mathrm{Diff}_{\nu}(\mathbb{A})$ corresponding to diagonal matrices in the matrix case. The “projection onto the diagonal” of a Lie algebra element is then given by averaging
$\pi\colon\mathfrak{a}\ni\varphi\mapsto\int_{-1}^{1}\varphi(\phi,\cdot)d\phi\in\mathfrak{d}=T_{e}\mathrm{D}.$
The corresponding “Weyl group” $\mathrm{WG}$ consists of maps $(\phi,z)\mapsto(j(z)\phi,a(z))$, where $a:[-1,1]\to[-1,1]$ is invertible, measure preserving, and $j\colon[-1,1]\to\{-1,1\}$ is measurable. In the smooth category, $\mathrm{WG}$ has only two elements, with $a(z)=z$ and $j(z)=\pm 1$. However, the idea of BFR is to work in a weaker topology, such that the Weyl group can “rearrange” the level-sets of functions in $\mathfrak{d}$ corresponding to changing the order of elements of diagonal matrices.
Let $\overline{\mathrm{Diff}_{\nu}(\mathbb{A})}$ denote the semigroup of measure preserving endomorphisms, which is the completion of $\mathrm{Diff}_{\nu}(\mathbb{A})$ with respect to the strong operator topology obtained by thinking of $\Phi\in\mathrm{Diff}_{\nu}(\mathbb{A})$ as the unitary operator on $L^{2}(\mathbb{A})$ given by $\omega\mapsto\omega\circ\Phi$. Let $\overline{\mathrm{WG}}$ denote the corresponding completion of the Weyl group, which does allow rearranging.
The first result of BFR is the convexity theorem, which is a beautiful, infinite-dimensional analog of Schur’s theorem for matrices.
Theorem 9 (Convexity theorem [5]).
Let $\zeta\in L^{2}(\mathbb{A})$ be a bounded, non-increasing, right continuous function of $z$ alone, and let $\overline{\mathcal{O}}(\zeta)=\zeta\circ\overline{\mathrm{Diff}_{\nu}(\mathbb{A})}$ be its orbit. Then $\pi(\overline{\mathcal{O}}(\zeta))\subset L^{2}([-1,1])$ is a weakly compact, convex set. Furthermore, the set of extreme points of $\pi(\overline{\mathcal{O}}(\zeta))$ is the orbit $\zeta\circ\overline{\mathrm{WG}}$.
The second result of BFR is the orbit theorem, which is a beautiful, infinite-dimensional analog of Horn’s theorem for matrices and the result that two (skew-)Hermitian matrices $A$ and $B$ share the same spectrum if and only if $\operatorname{tr}(A^{m})=\operatorname{tr}(B^{m})$ for $m=1,\ldots,N$.
Theorem 10 (Orbit theorem [5]).
If $\omega\in L^{\infty}(\mathbb{A})$, there is a unique bounded, non-increasing, and right continuous function $\zeta$ of $z$ alone such that $\omega\in\overline{\mathcal{O}}(\zeta)$. Furthermore, the orbit $\overline{\mathcal{O}}(\zeta)$ consists of all functions $\sigma\in L^{\infty}(\mathbb{A})$ with the same Casimirs as $\omega$, i.e.,
$\int_{\mathbb{A}}\sigma^{m}\,\nu=2\pi\int_{-1}^{1}\zeta^{m}\,d\phi,\quad m=1,2,\ldots.$
These two results are directly related to the question of completion of coadjoint orbits, flattened Reeb graphs, and vortex mixing as discussed above. Furthermore, the convexity theorem is directly related to mixing operators and the canonical splitting presented in Sec. 5 below.
We now translate the results of BFR to the spherical setting. Archimedes demonstrated that the horizontal projection of the cylinder to the sphere is area-preserving (i.e., symplectic). [^4] For this reason, V. Arnold called Archimedes “the first symplectic geometer”. Explicitly, it is the map $\alpha\colon\mathbb{A}\to\mathbb{S}^{2}$ given by
$\alpha\colon(\phi,z)\mapsto(\sqrt{1-z^{2}}\cos\phi,\sqrt{1-z^{2}}\sin\phi,z).$
It is almost everywhere invertible, with
$\alpha^{-1}\colon(x_{1},x_{2},x_{3})\mapsto(\arg(x_{1}+\mathrm{i}x_{2}),x_{3}).$
Notice that smooth functions in $\mathfrak{a}$ are constant exactly on the set-valued points
$\alpha^{-1}(0,0,\pm 1)=\big{(}\mathbb{S}^{1},\pm 1\big{)}.$
Since it is measure preserving $\alpha^{*}\mu=\nu$, it follows that $\overline{\mathrm{Diff}_{\nu}(\mathbb{A})}$ and $\overline{\mathrm{Diff}_{\mu}(\mathbb{S}^{2})}$ are isomorphic via
$\overline{\mathrm{Diff}_{\mu}(\mathbb{S}^{2})}=\alpha\circ\overline{\mathrm{Diff}_{\nu}(\mathbb{A})}\circ\alpha^{-1}.$
Furthermore, for $p\in\{1,\ldots,\infty\}$ any $f\in L^{p}(\mathbb{S}^{2})$ is mapped to $f\circ\alpha\in L^{p}(\mathbb{A})$ and vice versa. Thus, the convexity theorem and the orbit theorem translate to the spherical setting, such that “diagonal” functions are zonal (i.e., constant along longitudes).
The orbit theorem provides a natural closure of the smooth coadjoint orbits:
Proposition 11.
Let $\omega_{0}\in L^{\infty}(\mathbb{S}^{2})$ and let $\zeta\in L^{\infty}(\mathbb{S}^{2})$ be its corresponding unique non-increasing, right continuous zonal function (obtained via the orbit theorem). Then
$\overline{\mathcal{O}}(\omega_{0})\coloneqq\{\omega\in L^{\infty}(\mathbb{S}^{2})\mid\lambda_{\omega}=\lambda_{\omega_{0}}\}=\zeta\circ\overline{\mathrm{Diff}_{\mu}(\mathbb{S}^{2})}.$
Proof.
This result is essentially a reformulation of the orbit theorem. It follows since $\lambda_{\omega}=\lambda_{\omega_{0}}$ if and only if $\omega_{0}$ and $\omega$ have the same polynomial Casimirs (as already used in the proof of Theorem 6 above). ∎
The closure $\overline{\mathcal{O}}(\omega_{0})$ in Proposition 11 is located strictly in between the smooth coadjoint orbits $\mathcal{O}(\omega_{0})$ and Shnirelman’s weak closure $\overline{\mathcal{O}(\omega_{0})}^{*}$, i.e.,
$\mathcal{O}(\omega_{0})\subset\overline{\mathcal{O}}(\omega_{0})\subset\overline{\mathcal{O}(\omega_{0})}^{*}.$
In a sense, $\overline{\mathcal{O}}(\omega_{0})$ is the smallest closure that allows the matrix theory results to remain intact. It is also compatible with Yudovich’s global well-posedness theory, which gives a transport map $\Phi_{t}\in\overline{\mathrm{Diff}_{\mu}(\mathbb{S}^{2})}$ for initial data $\omega_{0}\in L^{\infty}(\mathbb{S}^{2})$.
Remark 6.
Bloch, Flaschka, and Ratiu [5] made a connection between matrices and area-preserving diffeomorphisms of the annulus via the Toda lattice model [55] and its continuum limit (the dispersionless Toda equations, cf. [6]). However, that model is somewhat unfulfilling, since the matrices are tridiagonal and do not “fill out” the infinite-dimensional group as they grow in size. [^5] The authors mention this and hint at a series corresponding to Hermitian matrices that “seem reasonable to consider”: they rediscovered quantization. Matrix hydrodynamics offers a remedy: a model that gives a direct link between the matrix groups $\mathrm{SU}(N)$ and area-preserving diffeomorphisms. In addition to Hamiltonian systems, the link also leads to matrix flow analogs of incompressible porous medium equations [33], via the gradient flow interpretation of the Toda lattice.
@@ 5. Canonical splitting and mixing operators
Stationary solutions of the Euler–Zeitlin equations (9) are characterized by the relation $[P,W]=0$. The stabilizer of $P$ is the subspace $\mathrm{stab}_{P}=\{W\in\mathfrak{su}(N)\mid[P,W]=0\}$. For a state $W$ and $P=-\Delta_{N}^{-1}W$, the distance of $W$ to $\mathrm{stab}_{P}$ thus measures how far from stationary the state is.
Let $\Pi_{P}$ denote the projection onto $\mathrm{stab}_{P}$ with respect to the inner product $\langle W,U\rangle_{2}=\frac{2\pi}{N}\operatorname{tr}(W^{\dagger}U)$ corresponding to $L^{2}$. It induces a canonical splitting $W=W_{s}+W_{r}$ where $W_{s}=\Pi_{P}W$. Explicitly, if $E\in\mathrm{U}(N)$ is an eigen basis for $P$, then $\Pi_{P}(W)=E^{*}\pi_{N}(EWE^{*})E$, where $\pi_{N}$ is the orthogonal projection onto the space of diagonal matrices (i.e., extracting the diagonal). From this formula, we can express the Euler–Zeitlin equations in the (now time dependent) variables $W_{s}$ and $W_{r}$:
@@
where $B$ is the unique element in the complement $\mathrm{stab}_{P}^{\perp}$ such that
$B,P]=\Pi_{P}^{\perp}\Delta^{-1}[P,W].$
(See Modin and Viviani [47, sect. 3.1] for a derivation of the equations (10).) By construction, the decomposition is orthogonal in $\langle\cdot,\cdot\rangle_{2}$. Curiously, it also fulfills a “reversed” orthogonally condition with respect to the energy norm $\lVert W\rVert_{E}^{2}\coloneqq\langle W,\Delta_{N}^{-1}W\rangle_{2}$.
@@
Remark 7.
In the special case when $P$ is diagonal, the projection $\Pi_{P}W$ selects the diagonal of $W$. From Schur’s theorem it then follows that $C_{f}^{N}(\Pi_{P}W)\leq C_{f}^{N}(W)$ for convex functions $f$, and that $C_{f}^{N}(\Pi_{P}W)=C_{f}^{N}(W)$ for all convex $f$ if and only if $P$ and $W$ are simultaneously diagonalizable. Since a general $P\in\mathfrak{u}(N)$ can be diagonalized, and since $(P,W)\mapsto\Pi_{P}W$ is equivariant under the adjoint action, we also have $C_{f}^{N}(\Pi_{P}W)\leq C_{f}^{N}(W)$ for generic $P$.
Why is the canonical splitting interesting? The theory of statistical mechanics, applied to Euler’s equations on $\mathbb{S}^{2}$, predicts a long-time behavior of small scale fluctuations about a large scale, steady state [26]. Thus, we expect $W_{r}=W-W_{s}$ to decrease [^6] But due to Poincaré recurrence only in a statistical sense: see zeitlin-reversibility-blog [44]. with time in the energy norm, which is in-sensitive to small scales. Indeed, in numerical simulations the component $W_{s}$ favors evolution into large scales, while $W_{r}$ favors small scales (see Fig. 5 and [47, Fig. 3b]). This canonical separation of scales is a perspective on Kraichnan’s [39] backward energy and forward enstrophy cascades.
@@
The analog of $\Pi_{P}$ is the operator $\Pi_{\psi}$ for $L^{2}$ projection onto $\mbox{stab}_{\psi}=\overline{\operatorname{ker}\{\psi,\cdot\}}^{L^{2}}$, formally given by averaging along the level-sets of $\psi$. For a generic $\psi\in C^{1}(\mathbb{S}^{2})$, it is well-defined as a norm-1 operator $\Pi_{\psi}\colon L^{p}(\mathbb{S}^{2})\to L^{p}(\mathbb{S}^{2})$ [47, prop. 5.4].
Remark 8.
Notice that the special case $\Pi_{x_{3}}$ is the projection $\pi$ in the Convexity theorem 9 of Bloch, Flaschka, and Ratiu [5]. If $\zeta$ is the zonal function from the Orbit theorem, so that $\psi=\zeta\circ\Phi$, and the measure preserving map $\Phi$ happens to be invertible, then we have the formula $\Pi_{\psi}(\omega)=\pi(\omega\circ\Phi^{-1})\circ\Phi$. From this perspective, we expect that $C_{f}(\Pi_{\psi}(\omega))=C_{f}(\omega)$ for all convex functions $f$ if and only if $\psi$ and $\omega$ are “simultaneously diagonalizable”, i.e., $\omega\circ\Phi^{-1}$ is zonal. This viewpoint is related to the work by Dolce and Drivas [16], which characterizes Shnirelman’s minimal flows via convex Casimirs.
Proposition 13 ([47]).
@@
Let $\{\omega_{n}\}$ be a uniformly bounded sequence in $L^{\infty}$. Then, for the limit $n\to\infty$, the following statements are equivalent:
  1. (1)
    $\|\omega_{n}-\overline{\omega}\|_{E}\rightarrow 0$,
  2. (2)
    $\omega_{n}\overset{\ast}{\rightharpoonup}\overline{\omega}$ weakly-* in $L^{\infty}$,
  3. (3)
    $\omega_{n}\rightharpoonup\overline{\omega}$ weakly in $L^{2}$.
@@
Weak-* convergence in $L^{\infty}$ implies weak convergence in $L^{2}$, since $L^{2}\subset L^{1}$, for bounded domains.
On the other hand, if $\omega_{n}\rightharpoonup\overline{\omega}$ weakly in $L^{2}$, then being $L^{2}$ compactly embedded in $H^{-1}$ by the Rellich–Kondrachov theorem, we have that $\|\omega_{n}-\overline{\omega}\|_{E}\rightarrow 0$.
Finally, it remains to prove that convergence in $H^{-1}$ implies convergence in the weak-* topology. Let $\phi\in C^{1}$, then we have that
$|\int(\omega_{n}-\overline{\omega})\phi dS|\leq\|\omega_{n}-\overline{\omega}\|_{E}\|\phi\|_{1}\rightarrow 0,\mbox{ for }n\rightarrow\infty,$
@@
for any $n\geq 1$. Hence, since $C^{1}$ is dense in $L^{1}$ and the sequence of linear functionals $f_{n}(\phi):=\int(\omega_{n}-\overline{\omega})\phi dS$ defined on $L^{1}$ is bounded in norm by the constant $C$ and so they are equicontinuous on $L^{1}$, we get that $\omega_{n}\overset{\ast}{\rightharpoonup}\overline{\omega}$, for $n\rightarrow\infty$. ∎
@@
The thesis follows from Lemma 15 and from the uniform bound
$\|\omega_{r}(t_{n})\|_{\infty}\leq 2\|\omega(0)\|_{\infty},$
for any $n\geq 1$. This bound is a consequence of [47, prop. 5.4], since
$\|\omega_{r}(t)\|_{\infty}=\|\omega(t)-\omega_{s}(t)\|_{\infty}\leq\|\omega(t)\|_{\infty}+\|\omega_{s}(t)\|_{\infty}\leq 2\|\omega_{0}\|_{\infty}.$
Hence, the states $\omega_{s}$, observed in our long-time numerical simulations, are natural candidates for the weak-* vorticity limits. We notice that minimal flows seem too restrictive as candidates for long-time asymptotics on $\mathbb{S}^{2}$; essentially, they allow only two vortex condensates, or the shape of similar signed condensates would have to balance perfectly, which is physically unlikely and numerically unseen. Instead, it is common to see branching in the functional dependence between vorticity and stream function [47, fig. 7 and 8].
@@
where $M$ is the angular momentum. Minimal flow with respect to these mixing operators might be compatible with the numerical observations (more than two blobs and branching), contrary to the full set of mixing operators, for which minimal flows cannot have branching and therefore generically necessitates two blobs only.